# Questioning the Survey Responses of Large Language Models

### v3

@@ 3.1 Testing for systematic biases: A-bias
It is well-known that language models’ most likely answer to multiple-choice questions can change depending on seemingly minor factors such as the ordering of few-shot examples [Zhao et al.,, 2021; Lu et al.,, 2022] or the ordering of answer choices [Robinson and Wingate,, 2023]. We are interested in the extent to which changes in choice ordering affect a model’s output distribution over answers.
@@
We measure A-bias for each question and model . Results are illustrated in Figure 3(b). We sort models by their size. We observe all models exhibit substantial A-bias. However, models in the order of a few billion parameters or fewer consistently exhibit particularly strong A-bias, and tend towards mono answers. We additionally observe that the strength of A-bias in instruction or RLHF tuned models is similar to that of base models, see Appendix B.2.
@@
We investigate other types of labelling and position bias (e.g., last-choice bias) in Appendix C. Overall, we find a strong tendency of LLMs to pick up on spurious signals in the way that answers are ordered and labeled, rather than their semantic meaning. Notably, in contrast to the primacy bias observed in humans [Groves et al.,, 2009], we find that models exhibit substantial A-bias even when randomizing the position of the “A” choice. Our findings are consistent with the concurrent work of Tjuatja et al., [2023], which similarly finds that models’ response biases to multiple-choice survey questions are generally not human-like.
@@
In summary, we find that systematic biases confound models’ answer distributions. This makes it challenging to draw robust conclusions about general properties of LLMs, and their comparison, from survey responses. For example, simply reversing the order of answers to the “SEX” question could lead to GPT-2 seemingly representing a population where females are significantly over-represented, whereas a reverse conclusion would be drawn when using the standard answer order. While much research went into designing the ACS to elicit faithful answers and eliminate systematic biases when surveying human populations, simply using the same question framing does not protect against the systematic response biases that language models exhibit.
4 Controlling for labeling and ordering biases
@@
To eliminate confounding due to labeling and ordering biases, we survey models under randomized choice ordering, borrowing an established methodology to adjust for ordering biases of all kinds in survey research [Groves et al.,, 2009]. In the following, we refer to the expected response after answer choice randomization as the adjusted response.
@@
In Figure 4(a) we plot the normalized entropy of base models’ adjusted responses for the ACS questions considered. We find that after adjustment, 1) the variations in responses’ entropy across survey questions are very small, 2) we no longer observe the trend of the entropy of model responses increasing log-linearly with model size. In fact, models’ survey responses have a normalized entropy of approximately irrespective of model size or survey question asked. This validates our initial hypothesis that, without adjustment, variations in responses across base models arise predominantly due to systematic biases such as A-bias, rather than the content of the survey questions asked.
@@ 4.1 Effect of instruction tuning
We now evaluate language models that have been fine-tuned with instructions and/or human preferences, henceforth “instruction-tuned models”. In Figure 4(b) we plot the normalized entropy of instruction-tuned models’ ACS survey responses after adjustment. We observe that instruction tuned-models all exhibit substantially higher variations in entropy across questions compared to base models. Nonetheless, the entropy of their responses is nonetheless higher than the entropy of the human responses observed in the U.S. census data.
@@
We now investigate the similarity of language models’ adjusted responses to the census data. To do so, we consider the overall U.S. census population, as well as 50 census subgroups corresponding to every state in the United States.
Inspired by the alignment measures proposed by Santurkar et al., [2023] and Durmus et al., [2023], we investigate the similarity of model responses to the census data by evaluating the average divergence across questions between model responses and the census statistics.333Whereas Santurkar et al., [2023] use the Wasserstein distance to compare answer distributions, we use KL divergence since questions in the ACS are predominantly nominal, rather than ordinal. We evaluate average KL divergence between each language model and each reference population , as follows:
@@
Looking at Figure 5(b) we find no consistent trend that instruction-tuning would move responses closer to the census, despite the increased deviation from uniform and the larger variations in entropy (recall Figure 4(b)). Only for larger models the divergence seems to clearly decrease with instruction-tuning. However, all models’ responses still remain significantly closer to the uniform baseline than to the U.S. census. For instance, the model with the largest number of responses being more similar to those of the U.S. census than to uniform is GPT-4, for which 6 out of 25 questions (24%) are closer to the U.S. census than to the uniform baseline.
4.3 Implications for survey-based alignment metrics
Our findings add important context to previous works studying the relative similarity between the survey responses of models and those of specific populations. In particular, due to the models’ tendency towards balanced answers, there is a strong correlation between a models’ similarity (or alignment) with a subgroup and the subgroup’s entropy, as shown in Figure 6. While we only depict the largest models in this figure, this correlation consistently holds for all models surveyed (see Appendix B.3). Interestingly, this trend also holds pre-adjustment, and we find that only the relative ordering of models changes with adjustment.
@@ 4.3 Implications for survey-based alignment metrics
Thus, our survey-derived alignment measure is more informative of the reference populations rather than the language models they aim to evaluate, since variations across subgroups tend to be more pronounced than variations across models. Particularities, such as the training data used or the demographics of the annotators used for fine-tuning with human feedback, seems to have little impact on which population is best represented. Models therefore consistently appear to be more “aligned” with the subpopulations exhibiting high entropy in their answers.
Beyond the ACS.
@@
To inspect whether this trend changes with the content of the questions asked, we reproduce our experiments with additional surveys. We use the American Trends Panel (ATP) opinion surveys considered by Santurkar et al., [2023], and the Pew Research’s Global Attitudes Surveys (GAS) and World Values Surveys (WVS) considered by Durmus et al., [2023]. These surveys encompass around 1500 questions and 60 U.S. demographic subgroups, and around 2300 questions and 60 national populations, respectively. We adopt the alignment metrics considered by the aforementioned works. We find that the insights gained from the ACS also hold for the ATP and GAS/WVS surveys.
@@
In particular, we similarly find a linear trend between the alignment metrics and subgroups’ entropy of responses, see Figure 18 in Appendix E.3. This observation explains some of the findings in prior works. For example, Santurkar et al., [2023] find that “all the base models share striking similarities–e.g., being most aligned with lower income, moderate, and Protestant or Roman Catholic groups” and “our analysis […] surfaces groups whose opinions are poorly reflected by current LLMs (e.g., 65+ and widowed individuals)”. For the ATP surveys considered, low income, moderate, and Protestant/Catholic are precisely the demographic subgroups with responses closest to uniformly random among the income, political ideology, and religion demographic subgroups; whereas age 65+ and widowed are the demographic subgroups with responses furthest from uniform among the age and marital status demographic subgroups.
5 Responses to entire questionnaires
In the previous sections, survey questions were presented independently of one another. We now seek to fill entire ACS questionnaires in a sequential manner, in order to generate for each language model a synthetic dataset of responses which emulates in form the ACS dataset collected by the U.S. Census Bureau. We then study the extent to which such synthetic datasets resemble the ACS dataset.
5.1 Methodology
We present survey questions in the same order as in the ACS questionnaire. When querying a model to answer survey question , we include a summary of the previously sampled answers in context. 444The maximum number of tokens in a filled questionnaire was less than 1024 tokens in all cases, thus fitting entirely within the context window of all surveyed models. We then sample from the model’s output probability distribution over answers, and continue to the next question. We illustrate this sequential process in Figure 7. We refer to Appendix D.3 for results collected with different variations of how a model’s previous answers are integrated into the prompt. We find our results to be robust to these prompt variations.
5.2 The discriminator test
5.3 Contrast with silicon samples
Argyle et al., [2022] propose “silicon sampling”, a methodology to produce synthetic survey respondents using LLMs by conditioning on actual survey respondents. They focus on a subset of 12 questions from the 2016 American National Election Studies (ANES) survey. For every human respondent, they construct a corresponding “silicon individual” by querying GPT-3 to predict the ANES respondent’s answer to each survey question given the respondent’s answers to all other questions. Their results indicate that, for the 2016 ANES survey, GPT-3 can be a fairly calibrated predictor of an individual’s answer to some survey question conditioned on the respondent’s answers to all other survey questions.555Lee et al., [2023] and Sanders et al., [2023] study imputation tasks similar to those of Argyle et al., [2022], and find that LLMs are not calibrated predictors for a variety of such tasks.
However, Argyle et al., [2022] emphasize that important insights can be gained by emulating the survey responses of human populations “prior to or in the absence of human data”. In this work we have considered precisely the setting where models’ responses are obtained in the absence of human data.666Conceptually, to generate each synthetic individual, we start with a blank survey questionnaire and prompt the LLM to sequentially fill the entire questionnaire. Whereas we only prompt LLMs with their own responses (i.e., to previous survey questions), Argyle et al., [2022] prompt the model only with actual human responses from the 2016 ANES data. To investigate how our findings transfer to the ANES, we reproduce the experiments of Section 5 using the 2016 ANES survey questionnaire considered by Argyle et al., [2022] and their “interview-style” prompt. We apply the discriminator test, and find that the trained classifiers can discriminate between the model-generated data and the ANES data with accuracy % (see Appendix E), indicating that models’ responses are markedly different to those in the ANES data.
@@ 6 Conclusion
We closely examined the survey responses of LLMs on the basis of the prime demographic survey conducted in the United States. To do so, we leveraged a popular methodology to elicit LLMs’ answer distributions to survey questions. We found that model responses are dominated by systematic ordering biases and do not exhibit the natural variations in entropy found in the human reference data collected by the US census. Even after adjusting for ordering biases, LLMs’ responses still do not resemble those of human populations. Instead, they exhibit consistently high entropy. This holds true both when presenting questions individually and in sequence, and irrespective of model size or fine-tuning with human preferences.
Taken together, our findings caution to expect robust insights when comparing LLMs’ responses against those of humans. In our study we could not find any indication that LLMs elicit faithful representations of any human population. It remains to be seen how to take advantage of the flexibility, scalability and scope of LLM survey data in the context of social science research. What seems clear is that LLMs’ survey data should be treated differently to human data and that the validity of surveys as a measurement instrument should be readdressed. The robustness and quality of an established survey does not seamlessly translate from the results obtained by surveying human populations to the logits output by LLMs.
@@ Appendix A Experimental details
We use the American Community Survey (ACS) Public Use Microdata Sample (PUMS) files made available by the U.S. Census Bureau.777https://www.census.gov/programs-surveys/acs/microdata.html The data itself is governed by the terms of use provided by the Census Bureau.888https://www.census.gov/data/developers/about/terms-of-service.html We download the data directly from the U.S. Census using the Folktables Python package [Ding et al.,, 2021]. We download the files corresponding to the year 2019.
@@
We open source the code to replicate all experiments.999https://github.com/socialfoundations/surveying-language-models In addition, the repository contains notebooks to visualize the results of our investigations under different prompt ablations.
@@ Appendix B Detailed experimental results
B.1 Model responses across questions before and after adjusting for A-bias
The results in this section complement Section 3, and pertain non-instruction-tuned language models. When surveying models without choice order randomization, we observe that the entropy of model responses tends to increase log-linearly with model size, often matching the entropy of the uniform distribution for the larger models. This trend is consistent across survey questions, irrespective of the question’s distribution over responses observed in the U.S. census (Figure 9).
B.2 A-bias of instruction-tuned models
The results in this section complement Section 3.1, and pertain instruction-tuned language models as well as language models fine-tuned with reinforcement learning with human feedback (RLHF). We observe that the strength of A-bias for these models, plotted in Figure 10, is comparable to that of base pre-trained models, plotted in Figure 3(b). This motivates the use of choice-order randomization in order to eliminate confounding due to labeling biases in models’ responses.
B.3 Relative alignment across demographic subgroups
The results presented here complement those of Section 4.3. We plot the average KL divergence between each language model and each demopgrahic subpopulation (U.S. state) against the average entropy of the subgroup’s responses. For readability, we split models into GPT-2 and GPT-Neo (Figure 11(a)), OpenAI’s API models (Figure 11(b)), MPT, Pythia, GPT-NeoX and its instruction variants (Figure 11(c)), and LLaMA, Llama 2 and its instruction and chat variants (Figure 11(d)).
@@ Appendix D Prompt ablations
We reproduce our experiments using different prompts to query the model. Due to the cost of querying OpenAI’s models, we only perform these ablations for models with publicly available weights. The notebooks with all figures can be retrieved from our Github repository.101010https://github.com/socialfoundations/surveying-language-models/blob/main/prompt-ablations Overall, the prompt ablation results are very consistent with the findings presented in the main text of the paper. In the following we provide an overview over the different ablations performed.
@@ Appendix E Results for ATP, GAS, WVS, and ANES surveys
We reproduce the experiments of Sections 3 and 4 using the ATP, and GAS/WVS used by Santurkar et al., [2023] and Durmus et al., [2023], where questions are presented individually of one another. We additionally reproduce the experiments of Section 5 using the 2016 ANES questionnaire considered by Argyle et al., [2022], where questions are presented in sequence. We do not consider OpenAI’s models as the cost to reproduce the experiments via the OpenAI API exceeds our budget. We obtain very similar results to those of the ACS presented in the main text of the paper. The notebooks with all figures can be retrieved from our Github repository.111111https://github.com/socialfoundations/surveying-language-models
@@ E.1 ATP surveys
We obtain the ATP survey questions and their corresponding human responses from the OpinionsQA repository.121212https://github.com/tatsu-lab/opinions_qa We present all answer choices when querying the models, but exclude the answer choices corresponding to refusals from our analysis similarly to Santurkar et al., [2023]. When comparing the similarity of models’ responses to different demographic subgroups, we use the demographic subgroups and the alignment metric considered by Santurkar et al., [2023]. For such metric, higher values of alignment indicate that models’ responses are more similar to the reference demographic group. We find that all models are more “aligned” with the uniformly random baseline than with any of the demographic subgroups, see Figure 16.

### v4

@@ 3.1 Testing for systematic biases: A-bias
It is well-known that language models’ most likely answer to multiple-choice questions can change depending on seemingly minor factors such as the ordering of few-shot examples [Zhao et al.,, 2021; Lu et al.,, 2022] or the ordering of answer choices [Robinson and Wingate, 2023a, ]. We are interested in the extent to which changes in choice ordering affect a model’s output distribution over answers.
@@
We measure A-bias for each question and model . Results are illustrated in Figure 3. We again sort models by their size. We observe all models exhibit substantial A-bias. However, models in the order of a few billion parameters or fewer consistently exhibit particularly strong A-bias, and tend towards mono answers. We additionally observe that the strength of A-bias in instruction or RLHF tuned models is similar to that of base models. A plausible explanation for small models exhibiting strong A-bias is that the ability to answer MMLU-style multiple-choice questions only emerges for models of sufficient scale [Dominguez-Olmedo et al.,, 2024].
@@
We investigate other types of labeling and position bias (e.g., last-choice bias) in Appendix C. Overall, we find a strong tendency of LLMs to pick up on spurious signals in the way that answers are ordered and labeled, rather than their semantic meaning. Notably, in contrast to the primacy bias observed in humans [Groves et al.,, 2009], we find that models exhibit substantial A-bias even when randomizing the position of the “A” choice. Our findings are consistent with the concurrent work of Tjuatja et al., [2023], which similarly finds that models’ response biases to multiple-choice survey questions are generally not human-like. Wang et al., [2024] additionally shows that models’ responses to multiple-choice survey questions may not consistently reflect their free-form outputs.
@@
In summary, we find that systematic biases confound models’ answer distributions. This makes it challenging to draw robust conclusions about inherent properties of LLMs, such as the opinions or populations they best represent. For example, simply reversing the order of answers to the “SEX” question could lead to GPT-2 seemingly representing a population where females are significantly over-represented, whereas a reverse conclusion would be drawn when using the standard answer order. While much research went into designing the ACS to elicit faithful answers and eliminate systematic biases when surveying human populations, simply using the same question framing does not protect against the systematic response biases that language models exhibit.
4 Inspecting adjusted responses
@@
To eliminate confounding due to labeling and ordering biases, we survey models under randomized choice ordering, borrowing an established methodology to adjust for ordering biases of all kinds in survey research [Groves et al.,, 2009]. Also a recent work in LLM research adopts this methodology [Robinson and Wingate, 2023b, ]. In the following, we refer to the expected response after answer choice randomization as the adjusted response.
@@
In Figure 4 we plot the normalized entropy of models’ adjusted responses for the ACS questions considered. First focusing on base models, and comparing the results to Figure 2(b) we find that after adjustment, 1) the variations in responses’ entropy across survey questions are very small, 2) we no longer observe the trend of the entropy of model responses increasing log-linearly with model size. In fact, models’ survey responses have a normalized entropy of approximately irrespective of model size or survey question asked. This validates our initial hypothesis that, without adjustment, variations in responses across base models arise predominantly due to systematic biases such as A-bias, rather than the content of the survey questions asked.
@@ 4.1 Effect of instruction tuning
We now evaluate language models that have been fine-tuned with instructions and/or human preferences, henceforth “instruction-tuned models”. In the right plot of Figure 4 we show the normalized entropy of instruction-tuned models’ ACS survey responses after adjustment. We observe that instruction tuned-models all exhibit substantially higher variations in entropy across questions compared to base models. But in general, the entropy of their responses remains higher than the entropy of the census responses. Interestingly, as we will see, although deviating more from uniform, model responses do not tend to be closer to the U.S. census responses.
@@
We now investigate the similarity of language models’ adjusted responses to the census data. To do so, we consider the overall U.S. census population, as well as 50 census subgroups corresponding to every state in the United States. This leads to different human reference populations.
Inspired by the alignment measures proposed by Santurkar et al., [2023] and Durmus et al., [2023], we investigate the similarity of model responses to the census data by evaluating the average divergence across questions between model responses and the census statistics.444Whereas Santurkar et al., [2023] use the Wasserstein distance to compare answer distributions, we use KL divergence since questions in the ACS are predominantly nominal, rather than ordinal. As we focus on categorical questions, we evaluate average KL divergence between each language model and each reference population , as follows:
@@
Looking at Figure 5 we find no consistent trend that instruction-tuning would move responses closer to the census, despite the increased deviation from uniform and the larger variations in entropy (recall Figure 4). Only for larger models the divergence seems to clearly decrease with instruction-tuning. However, all models’ responses still remain significantly closer to the uniform baseline than to the U.S. census. For instance, for the GPT-4 model whose answers exhibit the highest similarity to the human reference populations, only 6 out of 25 questions (24%) are closer to the U.S. census than to the uniform baseline. Given these results, drawing conclusions about the relative alignment of models with subgroups is prone to resulting in brittle conclusions.
5 Implications for survey-based alignment metrics
Our findings add important context to previous works studying the alignment of language models with different human subpopulations. In particular, we highlighted the tendency of models towards balanced answers. Due to varying entropy in the responses of subgroups this leads to a strong correlation between model alignment and the reference population’s entropy. The linear trend in Figure 6 visualizes this. For any given model, it consistently appears to be more “aligned” with the subpopulations exhibiting high entropy in their answers. Interestingly, we find that this trend also holds pre-adjustment, suggesting that the transformation of the response through randomized choice ordering is orthogonal to differentiating aspects of any specific population. In contrast, when comparing different models in Figure 6, we can see how adjustment has a large influence on their relative order. Differences across models that we see under naive prompting disappear after adjustment, which means that are largely attributable to systematic biases, rather than inherent properties of the models.
@@ 5 Implications for survey-based alignment metrics
Taken together our findings imply that the survey-derived alignment measure is more informative of differences in the reference populations rather than the language models is aims to evaluate. Model particularities, such as the pre-training data used, instruction tuning or the use of reinforcement learning with human feedback, seem to have little impact on which population is best represented.
5.1 Beyond the ACS
@@
To inspect whether this trend changes with the content of the questions asked, we reproduce our experiments with additional surveys. We use the American Trends Panel (ATP) opinion surveys considered by Santurkar et al., [2023], and the Pew Research’s Global Attitudes Surveys (GAS) and World Values Surveys (WVS) considered by Durmus et al., [2023]. These surveys encompass around 1500 questions and 60 U.S. demographic subgroups, and around 2300 questions and 60 national populations, respectively. We adopt the alignment metrics considered by the aforementioned works. We find that our insights gained from the ACS also hold for the ATP and GAS/WVS surveys. In particular, we similarly find a linear trend between the alignment metrics and subgroups’ entropy of responses, in particular after adjustment, see Figure 7.
@@
This observation partially explains some of the findings in prior works. For example, Santurkar et al., [2023] find that “all the base models share striking similarities–e.g., being most aligned with lower income, moderate, and Protestant or Roman Catholic groups” and “our analysis […] surfaces groups whose opinions are poorly reflected by current LLMs (e.g., 65+ and widowed individuals)”. For the ATP surveys considered, low income, moderate, and Protestant/Catholic are precisely the demographic subgroups with responses closest to uniformly random among the income, political ideology, and religion demographic subgroups; whereas age 65+ and widowed are the demographic subgroups with responses furthest from uniform among the age and marital status demographic subgroups. Further, Santurkar et al., [2023] observe that RLHF can result in a “substantial shift […] towards more liberal, educated, and wealthy [demographic groups]”. Our results suggest that this could be an artifact of systematic biases. For the ATP surveys, we observe three outliers for which its alignment before adjustment is not correlated with the entropy of subgroup’s responses: Llama 2 70B Chat and the two Llama 3 Instruct models. These are the models with largest pre-training compute considered. However, after adjustment, the alignment trends of Llama 2 70B Chat and the Llama 3 Instruct models are remarkably similar to that of their corresponding base models and all other LLMs.
@@ 6 Conclusion
We used a popular methodology to elicit LLMs’ answer distributions to survey questions and closely examined the responses on the basis of the prime US demographic survey. We found that model responses are dominated by systematic ordering biases and do not exhibit the natural variations in entropy found in the human reference data collected by the US census. Even after adjusting for ordering biases, LLMs’ responses still do not resemble those of human populations. Instead, they exhibit consistently high entropy, independent of the question asked. This holds true irrespective of model size or fine-tuning with human preferences.
These findings have important implications for insights gained from survey-derived alignment metrics. In particular, it explains why models of varying size all exhibit the same trend: they are most aligned with subgroups who happen to have balanced answers for the survey questions under consideration. For all models and surveys considered, alignment appears to be a proxy for the entropy of subgroups, rather than an inherent property of the model, or its training data.
We want to reiterate that our focus lies on questioning a popular methodology of eliciting survey responses from large language models using multiple choice prompting. At the example of this methodology our results highlight an important pitfall and suggest caution to expect robust insights when comparing such responses against those of human populations. The robustness and quality of an established survey does not seamlessly translate from the results obtained by surveying human populations to the logits output by LLMs. More research is needed to design methodologies for getting insights into the inherent biases of LLMs and the population they might represent. Here public surveys and their accompanying data offer exciting potential and the could play an important role as a benchmarking tool for systematic evaluations of LLMs, see [Cruz et al.,, 2024] as an example. Although the use of survey data for LLM research has recently gained popularity, it still remains a widely under explored data source.
Acknowledgements
The authors would like to thank Frauke Kreuter and the Social Data Science and AI Lab at Ludwig-Maximilians-Universität Munich for inspiring discussions on an earlier version of this manuscript. Celestine Mendler-Dünner acknowledges financial support from the Hector Foundation.
@@ Appendix A Experimental details
We use the American Community Survey (ACS) Public Use Microdata Sample (PUMS) files made available by the U.S. Census Bureau.555https://www.census.gov/programs-surveys/acs/microdata.html The data itself is governed by the terms of use provided by the Census Bureau.666https://www.census.gov/data/developers/about/terms-of-service.html We download the data directly from the U.S. Census using the Folktables Python package [Ding et al.,, 2021]. We download the files corresponding to the year 2019.
@@
We open source the code to replicate all experiments.777https://github.com/socialfoundations/surveying-language-models In addition, the repository contains notebooks to visualize the results of our investigations under different prompt ablations.
@@ Appendix B Detailed experimental results
The results in this section complement Section 3, and pertain non-instruction-tuned language models.
B.1 Model responses across questions before adjusting for A-bias
When surveying models without choice order randomization, we observe that the entropy of model responses tends to increase log-linearly with model size, often matching the entropy of the uniform distribution for the larger models. This trend is consistent across survey questions, irrespective of the question’s distribution over responses observed in the U.S. census, see Figure 8.
@@ Appendix D Prompt ablations
We reproduce our experiments using different prompts to query the model. Due to the cost of querying OpenAI’s models, we only perform these ablations for models with publicly available weights. The notebooks with all figures can be retrieved from our Github repository. 888https://github.com/socialfoundations/surveying-language-models/blob/main/prompt-ablations
Overall, the prompt ablation results are very consistent with the findings presented in the main text of the paper. In the following we provide an overview over the different ablations performed.
@@ Appendix E Results for ATP, GAS, WVS, and ANES surveys
We reproduce the experiments of Sections 3 and 4 using the ATP, and GAS/WVS used by Santurkar et al., [2023] and Durmus et al., [2023], where questions are presented individually of one another. We additionally reproduce the experiments of Section 5 using the 2016 ANES questionnaire considered by Argyle et al., [2023], where questions are presented in sequence. We do not consider OpenAI’s models as the cost to reproduce the experiments via the OpenAI API exceeds our budget. We obtain very similar results to those of the ACS presented in the main text of the paper. The notebooks with all figures can be retrieved from our Github repository. 999https://github.com/socialfoundations/surveying-language-models
@@ E.1 ATP surveys
We obtain the ATP survey questions and their corresponding human responses from the OpinionsQA repository.101010https://github.com/tatsu-lab/opinions_qa We present all answer choices when querying the models, but exclude the answer choices corresponding to refusals from our analysis similarly to Santurkar et al., [2023]. When comparing the similarity of models’ responses to different demographic subgroups, we use the demographic subgroups and the alignment metric considered by Santurkar et al., [2023]. For such metric, higher values of alignment indicate that models’ responses are more similar to the reference demographic group. We find that all models are more “aligned” with the uniformly random baseline than with any of the demographic subgroups, see Figure 13.